# Should You Use Your Large Language Model to Explore or Exploit?

### v1

@@ 1 Introduction
There has been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems. Early work in this direction has already produced impressive agentic behavior in both virtual [e.g.,  41, 32] and physical-world environments [e.g.,  7]. However, significant challenges remain.
Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making the best decision given the current data) and exploration (trying new options for long-term benefit). Balancing the two has been a subject of an enormous literature, see books [e.g.,  see books 38, 25, 3].
A recent line of work [e.g.,  23, 31] evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  by specifying the problem description, parameters, and history in the LLM prompt. Focused on simple tasks in reinforcement learning, these results have been mixed. Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so, e.g. by providing various summary statistics in-context, or by fine-tuning on data from algorithmic baselines on similar problem instances. Unfortunately, such mitigations do not readily extend beyond simple settings. For example, succinct summary statistics do not exist for many decision-making tasks, and fine-tuning for specific tasks may be prohibitive due to cost or lack of sufficient training data.
Motivated by these observations, we study the ability of Gpt-4 [2], Gpt-4o [18], and Gpt-3.5 [8] to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent. We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff.
@@
In Section 4, we evaluate LLMs as exploration oracles which suggest candidate actions within a large action space. To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (where the embedding is computed exogenously). Given the high dimensionality of the action space, traditional bandit approaches are inapplicable. However, an LLM can generate a small set of candidate actions which can then be used to instantiate an off-the-shelf bandit algorithm. We experiment with several prompting strategies, and find that they all lead to relatively good exploration. Finally, we repeat our experiments on a larger-scale bandit task based on paper titles and abstracts from arXiv (where the goal is to find a suitable title for a given abstract), with similar findings.
2 Related Work and Background
@@
Our results belong to a small, but quickly growing line of work on using pre-trained LLMs for in-context reinforcement learning (RL). Coda-Forno et al. [13], Krishnamurthy et al. [23], Nie et al. [31], Monea et al. [30], Xia et al. [43], Park et al. [33], Wu et al. [42] evaluate the ability of LLMs to solve various multi-armed bandit and contextual bandit tasks, and find that the current generation of LLMs largely fail to solve these tasks in-context. Indeed, positive findings are restricted to very simple tasks and/or require substantial mitigations (which in turn do not readily extend beyond simple settings). Xia et al. [43] use LLMs to solve dueling bandit tasks, and Park et al. [33] also evaluate the ability of LLMs to learn in games. While our paper is primarily concerned with whether LLMs succeed as algorithms, several others [e.g.,  36, 17, 14] use in-context bandits (and many other tasks) to study whether LLMs exhibit human-like behavior/biases in decision-making.
@@
A broader literature on in-context learning [8] aims to solve various tasks by providing all relevant information in the LLM prompt. The work on exemplar selection (selecting examples and other information to present in-context) [e.g.,  19, 48, 44, 40] is relevant to our exploitation experiments.
@@
A growing line of work aims to use LLMs as a part of a larger decision-making agent [e.g.,  28, 50, 49]. Our exploration experiments take inspiration from the work on using LLMs as “action priors” inside of a larger RL algorithm [46, 10, 47, 16]. Much of this work falls under the proposer-verifier framework of Snell et al. [39], where an LLM proposes several possible sequences from which a verifier selects suitable candidates. In comparison, our goal is a more systematic evaluation of LLMs’ abilities to explore large action spaces, in isolation from other components of the decision-making task.
Finally, a parallel line of work trains transformers from scratch to solve various RL tasks [e.g.,  24, 26, 34, 45, 27].
@@
Multi-armed bandits. We consider tasks based on multi-armed bandits (MAB) and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see books [38, 25] for background. In MAB, there are $T$ rounds and $K$ arms. In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes a noisy reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm. The reward distribution, and particularly its mean $\mu(a_{t})$, are unknown to the algorithm. In CB, the learner additionally observes a context $z_{t}$ before each round $t$, and the expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm. The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
@@
An “exploitation oracle" (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader. Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification [38, Ch.8]. Designing CB exploitation oracles for their own sake, a.k.a. offline policy optimization, is well-studied [starting from, e.g.,  6, 15].
An “exploration oracle" can suggest a set of candidate actions to discretize a large action space and initialize an off-the-shelf bandit algorithm. Ideally, the oracle would "zoom in" on a lower-dimensional subspace with better-performing actions. Our intuition for exploration oracles is grounded in the study of bandits with Lipschitz-like structure [38, Ch.4]. In particular, the size vs. quality tradeoff in the discretization and the importance of “zooming in" are key issues in this literature [e.g.,  22, 20, 21, 9].
@@ 3 LLMs as Exploitation Oracles
In this section, we evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes. We present LLMs with in-context exploit tasks inspired by multi-armed bandits (MAB) and contextual bandits (CB). In a CB exploit task, an LLM is given a history consisting of context-arm-reward tuples, and is instructed to take the best arm given the current history and the current context. A MAB exploit task is the same, but without contexts. These tasks are generated from some parameterized distributions called exploit puzzles. Our visualizations plot the fraction of correct answers against the task difficulty.
3.1 MAB Exploit Puzzles
Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with "buttons" prompt. The following conventions apply to all figures in this section. Each line corresponds to a particular value of $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis. The shaded band around the line represents a $95\%$ confidence interval. The dashed line is the number of tasks (“runs") with empirical gap at most $\epsilon$; the corresponding Y-scale is on the right.
@@ 3.1 MAB Exploit Puzzles
Once an exploit task is generated, we measure its difficulty via empirical gap: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$. Intuitively, the smaller its empirical gap, the more difficult is the puzzle. Let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
@@
We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer": an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$. We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level. Hence, we plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against empirical gap $\epsilon$.
@@
We find that Gpt-4 and Gpt-3.5 do not perform well on these MAB exploit puzzles, see Figure 1. Performance tends to degrade (1) as the history length $T$ increases and (2) as the empirical gap decreases. While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4. We found that performance was similar across all of our prompt designs (see Appendix A for more plots).
3.2 CB Exploit Puzzles and Mitigations
@@
Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left) and fails on a slightly larger one (right).
@@
While the history in $K$-armed bandits can be summarized using $2K$ numbers (the average reward for each arm and number of times it has been played), such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB. [^2] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
We focus on linear CB, a well-studied CB model [starting from 29, 12, 1] in which the expected reward of each arm is linear in the context. Specifically, given context $z\in\mathbb{R}^{d}$ and arm $a$, the expected reward is $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$, for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.
We consider a CB exploit puzzle parameterized by the number of arms $K$, context dimension $d$, and history size $T$. The tasks are constructed as follows. We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-0.25,0.25]$ independently and u.a.r. for each arm $a$. Given context $z\in\mathbb{R}^{d}$, expected reward for arm $a$ is $\mu(z,a)=\langle z,\theta_{a}\rangle+\gamma_{a}$. We generate a history of $T$ rounds $t\in[T]$. Contexts $z_{t}$ are sampled independently and u.a.r. from $[-1,1]^{d}$. For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward is generated as $\mu(z_{t},a)$ plus an independent unit-variance Gaussian. Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action which appears best. This gives one exploit task. We generate $N$ tasks for the same $K,T,d$.
We present these exploit tasks to the LLM in-context using a modified “buttons” prompt, where contexts corresponds to “numbers on a screen" which affect the payoffs for pressing each button. The prompt does not mention linearity of the CB instance (because such model-based information is typically unavailable in applications).
For a particular exploit task, we now define a “correct answer" as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.  [^3] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context. As before, $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks.
Likewise, the task difficulty is not easily defined in terms of the realized rewards. Instead, we focus on the effective gap: the difference in expected reward between the best and second-best arm given the current context $z=z_{T+1}$. That is, the difference between the largest and second-largest number among $\mu(a,z)$, $a\in[K]$. Intuitively, smaller effective gap corresponds to increased difficulty.
[…]
Motivated by these observations, we implement several natural mitigations inspired by the literature on exemplar selection for in-context learning (discussed in Related Work). The mitigations are as follows:
@@
In all three mitigations, we do not explain the “mitigation strategy" in the prompt. That is, we present the reported context-arm-reward tuples as if it were the entire history, not mentioning nearest neighbors, clustering, or averaging.
Figure 3 visualizes the performance of these mitigations on a slightly larger (but still relatively small) puzzle with $K=d=2$ and $T=4000$. We use Gpt-4 with and without CoT. We compare the mitigations against linear regression baseline (which is effectively an "upper bound", as the underlying CB instance is linear). Without CoT prompting, we find that all three mitigations achieve $\mathrm{FracCorrect}$ around $80\%-90\%$, although this dips to around $60\%-85\%$ when using CoT. In addition to (potentially) improving performance, mitigations can also offer a practical way to solve decision-making tasks using LLMs when the history is large; when $T=4,000$, our prompt vastly exceeds the context window of all models we had access to.
Figure 5: Gpt-4o on the text-based CB exploit puzzle: some mitigations help, but are outperformed by linear regression.
@@
3.3 CB Exploit Puzzles (text-based)
As a robustness check, we repeat our CB experiments on a text-based exploit puzzle. In this puzzle, contexts are items in a room (e.g. animals, objects on a table), and actions have an associated semantic meaning (e.g. eat the food item, leave the room). Rewards are still presented numerically, and are non-linear functions of both the context and action. See Appendix A for full details on our experimental setup.
Figure 5 shows the performance of Gpt-4o (with mitigations) on this puzzle. While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline.

### v2

@@ 1 Introduction
There’s been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems. Early work in this direction has already produced impressive agentic behavior in both virtual (e.g.,  Wang et al., 2023; OpenAI, 2025) and physical-world environments (e.g.,  Black et al., 2024).
Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making the best decision given the current data) and exploration (trying new options for long-term benefit). Balancing the two has led to an enormous literature (e.g.,  see books Slivkins, 2019; Lattimore and Szepesvári, 2020; Agarwal et al., 2020).
A recent line of work (e.g.,  Krishnamurthy et al., 2024; Nie et al., 2024) evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  specifying the problem description, parameters, and history in the LLM prompt. Focused on simple tasks in reinforcement learning (RL), these results are mixed. Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so by providing succinct summary statistics in-context. However, such statistics do not exist beyond simple decision-making problems, e.g.,  for contextual bandits. RL-specific pre-training or fine-tuning (on data from algorithmic baselines on similar problem instances) tends to work well (see Related Work), but may be prohibitive due to cost or insufficient training data. Besides, using a commonly available frontier LLM would leverage its “generalist” intelligence and may be much easier in terms of logistics and required expertise.
Motivated by these observations, we study the ability of LLMs to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent. We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff. We experiment with a variety of LLMs: Gpt-4, Gpt-4o, Gpt-3.5, Qwen-2.5, Gemma-3, Mistral-7B, as well as DeepSeek-R1, a reasoning model.  [^2] Several other LLMs did not work for our purposes: Llama-3.2-3B and Llama-3-8B would not follow instructions, as well as DeepSeek-R1-Distill-Llama-70B (a reasoning model); Qwen-2.5-Math-70B (another reasoning model that we tried) has a short context window. We restrict our GPT access to Gpt-4, Gpt-4o, and Gpt-3.5, as newer variants such as Gpt-4.1 and Gpt-4.5 are primarily optimized for efficiency (e.g. latency, cost, coding) rather than broad general reasoning, and Gpt-5 uses a reasoning backbone as a routing mechanism across sub-models.
@@
In Section 3, we evaluate LLMs as exploration oracles which suggest candidate actions by discretizing a large action space. To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (the latter is computed exogenously). Given the high dimensionality of the action space, traditional discretization approaches typically used in continuous bandit problems are inapplicable. However, an LLM can generate a small set of candidate actions which can then be used to instantiate an off-the-shelf bandit algorithm. We experiment with several prompting strategies, and find that they all lead to relatively good exploration. Finally, we repeat our experiments on a larger-scale bandit task based on paper titles and abstracts from arXiv (where the goal is to find a suitable title for a given abstract), with similar findings.
@@
Related work. Our results belong to a growing line of work on using pre-trained LLMs for in-context reinforcement learning (RL). Coda-Forno et al. (2023); Krishnamurthy et al. (2024); Nie et al. (2024); Monea et al. (2024); Xia et al. (2024); Park et al. (2024); Wu et al. (2024) evaluate the ability of LLMs to solve various multi-armed bandit and contextual bandit tasks, and find that the current generation of LLMs largely fail to solve these tasks in-context. Indeed, positive findings are restricted to very simple tasks and/or require substantial mitigations (which in turn do not readily extend beyond simple settings). Xia et al. (2024) use LLMs to solve dueling bandit tasks, and Park et al. (2024) also evaluate the ability of LLMs to learn in games. While our paper is primarily concerned with whether LLMs succeed as algorithms, several others (e.g.,  Schubert et al., 2024; Hayes et al., 2024; Coda-Forno et al., 2024) use in-context bandits (and many other tasks) to study whether LLMs exhibit human-like behavior/biases in decision-making.
@@
A broader literature on in-context learning (starting from Brown et al., 2020) aims to solve various tasks by providing all relevant information in the LLM prompt. The work on exemplar selection (selecting examples and other information to present in-context) (e.g.,  Khalifa et al., 2023; Zhang et al., 2022; Xiong et al., 2023; Tonglet et al., 2023) is relevant to our exploitation experiments.
@@
A growing line of work aims to use LLMs as a part of a larger decision-making agent (e.g.,  Li et al., 2024; Zhou et al., 2023; Zhao et al., 2024). Our exploration experiments take inspiration from the work on using LLMs as “action priors” inside of a larger RL algorithm (Yan et al., 2024; Carta et al., 2023; Yao et al., 2024; Hao et al., 2023). Much of this work falls under the proposer-verifier framework of Snell et al. (2024), where an LLM proposes several possible sequences from which a verifier selects suitable candidates. In comparison, our goal is a more systematic evaluation of LLMs’ abilities to explore large action spaces, in isolation from other components of the decision-making task.
Finally, a parallel line of work trains transformers to solve various RL tasks (e.g.,  Laskin et al., 2022; Lin et al., 2023; Lee et al., 2024; Raparthy et al., 2023; Xu et al., 2022; Lehnert et al., 2024; Mukherjee et al., 2024). Our negative results in Section 2 provide an additional foundation for this work, emphasizing the shortcomings on frontier LLMs that are not custom-trained for RL.
@@
Background on multi-armed bandits (MAB). We consider tasks based on MAB and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see Slivkins (2019); Lattimore and Szepesvári (2020) for background. In MAB, there are $T$ rounds and $K$ arms. In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm. The reward distribution, and particularly its mean $\mu(a_{t})$, are unknown to the algorithm. In CB, the learner additionally observes a context $z_{t}$ before each round $t$, and the expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm. The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
@@
An “exploitation oracle” (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader. Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification (Slivkins, 2019, Ch.8). Designing CB exploitation oracles for their own sake, a.k.a. offline policy optimization, is well-studied (starting from, e.g.,  Beygelzimer and Langford, 2009; Dudík et al., 2014). Usage of an LLM as an “exploration oracle” is closely related to the literature on dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011), as we elaborate in Section 3.
@@ 2 LLMs as exploitation oracles
Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with ”buttons” prompt. The following conventions apply to all figures in this section. Each line corresponds to a particular value of #rounds $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis. The shaded band around the line represents a $95\%$ confidence interval. The dashed line is the number of tasks (“runs”) with empirical gap $\leq\epsilon$; the resp. Y-scale is on the right.
We evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes. We present LLMs with in-context exploit tasks inspired by multi-armed bandits (MAB) and contextual bandits (CB). In a CB exploit task, an LLM is given a history consisting of context-arm-reward tuples, and is instructed to take the best arm given the current history and the current context. A MAB exploit task is the same, but without contexts. These tasks are generated from some parameterized distributions called exploit puzzles.
As our results here are largely negative, we focus on relatively small tasks. We use synthetic data, to better control task size/difficulty and trace the “boundary of tractability”. We find that LLMs are consistently outperformed by a simple linear regression baseline, even on moderately-sized tasks. Due to computational constraints, our experiments focused on Gpt-4, Gpt-4o, and Gpt-3.5, with additional LLMs evaluated as a robustness check in Appendix A.4, with similar findings.
MAB exploit puzzles. Our MAB-based experiments on Gpt-4 and Gpt-3.5 provide a partial explanation for why the current LLMs fail to solve MAB tasks in-context when presented with raw (non-summarized) history, as first observed by Krishnamurthy et al. (2024); Nie et al. (2024). Following these two papers, we try two prompts: one in which arms correspond to pushing different colored buttons and one where they correspond to showing different advertisements to users. The LLM is asked to choose the arm with the highest empirical reward in the next round. We also try chain-of-thought (CoT) prompts, for the total of 4 prompt designs: $\left\{\,\text{buttons, adverts}\,\right\}\times\left\{\,\text{CoT, no-CoT}\,\right\}$. See Appendix A for more details on our experimental setup.
@@
Given a (realized) exploit task, we measure its difficulty via empirical gap $\epsilon$: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$. Intuitively, puzzle difficulty decreases with $\epsilon$ (as one needs to distinguish the top two arms). Empirical gap mirrors the “gap” between the top two expected rewards, a standard notion of difficulty in MAB. Let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
@@
We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer”: an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$. We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level. Hence, we plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against empirical gap $\epsilon$.
@@
We find that Gpt-4 and Gpt-3.5 do not perform well on these MAB exploit puzzles, see Figure 1. Performance tends to degrade (1) as the history length $T$ increases and (2) as the empirical gap decreases. (Our intuition for the former is that the LLM has more data points to process in context, and for the latter is that the bandit instance becomes more difficult.) While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4.333From CoT logs, we found that Gpt-4 was doing incorrect reasoning in a non-trivial fraction of runs. We found that performance was similar across all of our prompt designs (see Appendix A for more plots).
@@
Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left), but fails on a slightly larger one (right).
CB exploit puzzles and mitigations. While the history in $K$-armed bandits can be summarized using $2K$ numbers (for each arm, the average reward and #plays) such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB. [^4] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
We focus on linear CB, where the expected reward of each arm $a$ is linear in the context $z\in\mathbb{R}^{d}$: $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$ for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.  [^5] Linear CB are well-studied, starting from Li et al. (2010); Chu et al. (2011); Abbasi-Yadkori et al. (2011). (We try non-linear CB, too.)
We consider a CB exploit puzzle parameterized by #arms $K$, dimension $d$, and history size $T$. The tasks are constructed as follows. We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-0.25,0.25]$ independently and u.a.r. for each arm $a$. Given context $z\in\mathbb{R}^{d}$, expected reward for arm $a$ is $\mu(z,a)=\langle z,\theta_{a}\rangle+\gamma_{a}$. We generate a history of $T$ rounds $t\in[T]$. Contexts $z_{t}$ are sampled independently and u.a.r. from $1,1]^{d}$. For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward equals $\mu(z_{t},a)$ plus an independent unit-variance Gaussian. Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action for round $T+1$ which appears best. This gives one exploit task. We generate $N$ tasks for the same $K,T,d$.
We use a modified “buttons” prompt, where contexts correspond to “numbers on a screen” which affect the payoffs for each button. The prompt does not mention linearity of the CB instance (because such model-based information is typically unavailable in applications).
Given an exploit task, we now define a “correct answer” as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.  [^6] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context. $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks. Likewise, the task difficulty is not easily defined in terms of the realized rewards. Instead, we focus on the effective gap: the difference in expected reward between the best and second-best arm given the current context $z=z_{T+1}$. That is, the difference between the largest and second-largest number among $\mu(a,z)$, $a\in[K]$. Intuitively, smaller effective gap corresponds to increased difficulty.
First, we find that Gpt-4 obtains near-perfect performance on a “small” puzzle with $K=2$ arms, context dimension $d=1$, and history size $T=50$ (Figure 2). However, its performance degrades as the problem size increases: see Figure 2 for $K=d=2$ and $T=100$. Moreover, limited prompt size may prevent processing larger histories. [^7] E.g.,  our LLM access points bottomed out at $T\approx\text{100-200}$ for Gpt-4 and $T\approx\text{1000-2000}$ for Gpt-4o.
@@
Motivated by these observations, we implement several natural mitigations inspired by the literature on exemplar selection for in-context learning (discussed in Related Work).
@@
We do not explain the “mitigation strategy” in the prompt: we present the reported context-arm-reward tuples as if it were the entire history, without mentioning clustering or averaging.
Figure 3 visualizes the performance of these mitigations on a slightly larger (but still relatively small) puzzle with $K=d=2$ and $T=4000$. We use Gpt-4 with and without CoT. We compare the mitigations against linear regression baseline (which is effectively an ”upper bound”, as the underlying CB instance is linear). Without CoT prompting, we find that all three mitigations achieve $\mathrm{FracCorrect}$ around $80\%-90\%$, although this dips to around $60\%-85\%$ when using CoT. In addition to (potentially) improving performance, mitigations can also offer a practical way to solve decision-making tasks using LLMs when the history is large; when $T=4,000$, our prompt vastly exceeds the context window of all models we had access to.
@@
Figure 5: Left: Performance of Deepseek-R1 on our numerical CB puzzle. Right: Gpt-4o on the text-based CB exploit puzzle. Some mitigations help, but are outperformed by linear regression.
Reasoning models are LLMs specifically trained to follow multi-step chains of thought to solve complex problems that require intermediate reasoning. We ran one such model, DeepSeek-R1-Distill-Qwen-32B (henceforth DeepSeek-R1), a 32B parameter reasoning model distilled from Qwen-2.5. Our results are shown in Figure 5 (left). As before, we find that DeepSeek-R1 fails to match the performance of the baseline, even when using the mitigations. High inference costs and small context windows prevented us from running larger reasoning models.
CB exploit puzzles (text-based & non-linear). As a robustness check, we repeat our CB experiments on a text-based exploit puzzle. In this puzzle, contexts are items in a room (e.g. animals, objects on a table), and actions have an associated semantic meaning (e.g. eat the food item, leave the room). Rewards are still presented numerically, and are non-linear functions of both the context and action. See Appendix A for full details on our experimental setup.
Figure 5 (right) shows the performance of Gpt-4o (with mitigations) on this puzzle. While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline.
Our intuition for these findings is as follows: While LLMs are generally good with textual inputs, the reward data is numerical, so the LLM faces most/all of the same challenges as before (i.e.,  choosing a “solution strategy” and executing it correctly), exacerbated by non-linearity. Meanwhile, linear regression is known to often work fairly well even on non-linear data.
Takeaways. While unmitigated performance drops as $d,K,T$ grow, increases in $T$ can largely be taken care of by our mitigations, whereas increases in $d$ and $K$ cannot. Indeed, taken together, our findings in this section suggest that while the current generation of LLMs are effective at exploiting in small-scale decision-making tasks, they struggle in larger settings.